13.6 b. Find a normal vector to the level curve corresponding to c = In 5 for the function given by f(x,y)= In(x4 + y^) at thPlease do clearly and explain your steps!
Answer #1

Given f(x,y) = ln (x+y). c=lu s Point (x, y) = (-2,1). f(e,y) = c at P... The usual vector to the level one is the Sadient offg (1,4) - Sony (10 (2-0) Apply differentiation sale Cu Cuco) in wom). 24 fy(x,y) - 2 Jf (014) = fx i+fyi 0f68,5) = ang teles

Normal vector to the level curve is (-4/5) i+(2/5) j